By the method of separation of variables, solve the boundary value problem

∂u/∂x=4∂u/∂y, u(0,y)=8e-3y

1
Expert's answer
2021-11-05T13:51:00-0400

let the solution is of the form u(x,y)=X(x)Y(y).

put u(x,y) in the given pde.

"Y\frac{\partial X}{\partial x}=4X\frac{\partial Y}{\partial y}\\
\frac{1}{X}\frac{\partial X}{\partial x}=\frac{4}{Y}\frac{\partial Y}{\partial y}=\lambda(say)\\
\text{first solve}\\
\frac{1}{X}\frac{\partial X}{\partial x}=\lambda\\
\implies X=c_1e^{\lambda x}\\
\text{Second solve}\\
\frac{4}{Y}\frac{\partial Y}{\partial y}=\lambda\\
\implies Y=c_1e^{ \frac{\lambda y}{4}}\\
\text{Therefore, the solution is given by}\\
u(x,y)=c_1c_2e^{\lambda(x+\frac{ y}{4})}
=ce^{\lambda(x+\frac{ y}{4})}\\
\text{Now, apply the given condition, we get}\\
c=8e^{-y(3+ \frac{\lambda}{4})}
\text{Therefore, the solution is}\\
u(x,y)=8e^{-y(3+ \frac{\lambda}{4})}e^{\lambda(x+y)}\\
\text{ Since, it is a boundary value problem. Therefore, when the boundary conditions given then we can calculate the value of lambda and get the solution u(x,y).}"